Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability tea

Expert replies
by jzebra10 » Tue Nov 29, 2011 12:09 am
I don't understand the explanation given in the book. Can someone try to explain this better?

At a blind taste competition a contestant is offered 3 cups of each of the 3 samples of tea in a random arrangement of 9 marked cups. If each contestant tastes 4 different cups of tea, what is the probability that a contestant does not taste all of the samples?

a) 1/12
b) 5/14
c) 4/9
d) 1/2
e) 2/3

thanks!
Join the discussion
Source: — Problem Solving |

by Anurag@Gurome » Tue Nov 29, 2011 12:49 am
jzebra10 wrote:I don't understand the explanation given in the book. Can someone try to explain this better?

At a blind taste competition a contestant is offered 3 cups of each of the 3 samples of tea in a random arrangement of 9 marked cups. If each contestant tastes 4 different cups of tea, what is the probability that a contestant does not taste all of the samples?

a) 1/12
b) 5/14
c) 4/9
d) 1/2
e) 2/3

thanks!
Probability that a contestant does not taste all of the samples = 1 - probability that a contestant tastes all 3 samples.

This means that first we should find the probability that a contestant should taste 2 cups of one sample and 1 cup from each of 2 other samples.

Total no. of ways to choose 4 cups out of 9 = 9C4
No. of ways to choose the sample which provides with 2 cups = 3C1
No. of ways to choose these 2 cups from the sample chosen = 3C2
No. of ways to choose 1 cup out of 3 from the 2nd sample = 3C1
No. of ways to choose 1 cup out of 3 from the 3rd sample = 3C1

So, probability that a contestant tastes all 3 samples = (3C1 * 3C2 * 3C1 * 3C1)/9C4 = (3 * 3 * 3 * 3)/(18 * 7) = 9/14

Therefore, required probability = 1 - 9/14 = [spoiler]5/14[/spoiler]

The correct answer is B.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by rijul007 » Tue Nov 29, 2011 12:55 am
Total no of ways of tasting 4 cups of tea = 9P4 = 6*7*8*9

No of ways that all three samples are selected =>
Selecting two of a kind = 3*3C2 = 9
Other two cups different types = 3*3 = 9
Arranging them = 4!

Total no of ways = 9*9*4!

Probability that a contestant tastes all samples = P1 = 9*9*4!/6*7*8*9 = 9/14

probability that a contestant does not taste all of the samples = 1-P1 = 1- 9/14 = 5/14

IMO: B
Join the discussion

by shankar.ashwin » Tue Nov 29, 2011 12:58 am
You could also do it directly,

Since there are only 3 cups of each type, we can be sure he tastes 2 varieties of tea (Since he tastes a total of 4).

We have 3 samples and we taste only 2 - 3C2 = 3 ways of choosing which 2 to taste

And once we select the 2 samples, we have 6 cups of the 2 samples and we pick 4 - 6C4 = 15

Total ways = 9C4 = 126

So, we have (3*15)/126 = 5/14 (3C2*6C4/9C4)
Join the discussion

by HSPA » Tue Nov 29, 2011 12:58 am
Base: 4 out of 9 has to be chosen = 9C4

Does not taste all flavours = 1- tastes all flavours

Conditions to taste all flavours = 3*( 3c1*3c1*3c2) = 3^4

result = 1- 3^4/ 9c4 = 1- 9/14 = 5/14
First take: 640 (50M, 27V) - RC needs 300% improvement
Second take: coming soon..
Regards,
HSPA.
Join the discussion